one number is 3 times a first number. a third number is 100 more than the first number. if the sum of the three numbers is 585, find the numbers
Answer by Edwin McCravy(20081)

One number Y is 3 times a first number X. 

Y = 3X

A third number Z is 100 more than the first number X.

Z = X + 100 

The sum of the three numbers is 585

X + Y + Z = 585

Can you solve this system of equation:

ìY = 3X
íZ = X + 100
îX + Y + Z = 585

If not post again asking how.  The solution is

X = 97, Y = 291, Z = 197

Checking in the words, not the equations:

291 is 3 times 97.  197 is 100 more than the 97. 
The sum of 97, 291 and 197 is 585.  That checks.

Edwin